OpenStudy (anonymous):

Help Fast Prove: tan2 θ•cos2 θ + cos2 θ = 1.

OpenStudy (anonymous):

Please show step by step

OpenStudy (solomonzelman):

tan^2x times cos^2x = sin^2x b/c tan=sin/cos and the cosines cancel

OpenStudy (solomonzelman):

Then apply sin²x+cos²x=1

OpenStudy (solomonzelman):

or unless it is not ², but rather (2 theta) ... Am I getting your question right Miss ?

OpenStudy (anonymous):

Yes it's squared theta

OpenStudy (anonymous):

thank you so much for the help!

OpenStudy (solomonzelman):

Anytime !